Contents

General Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Commands for Power Users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Special Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Hint Book . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Perdition’s Flames

Copyright \begin{matrix}c \\ ◯ \end{matrix}1993 by Michael J. Roberts. All Rights Reserved.

Copyright and Software License Information

The Perdition’s Flames software and this manual are copyrighted works. You may use and distribute this software free of charge, subject to certain limitations. In particular, you must include this copyright information with all copies you make; you may not require or collect a fee for copies of this software that you give to other people; you may not include this software with any other software for which a fee is collected; you may not modify this software in any way, and each copy you make and distribute must be a full and complete copy of the software you originally received. Anyone to whom you give a copy of this software receives all of the same permissions that you did under this license.

Introduction to “Perdition’s Flames”

If you’ve ever played a text adventure game, you’ve certainly seen this dreaded message—probably at times you never expected it:

  *** You have died. ***

At this point, most games end, and you have to start over or go back to a saved position. Now, for the first time, a game takes you beyond “You have died” and sends you on a strange and challenging journey through the nether regions.

If you’ve played other adventure games, you’ll probably be surprised (and pleased, we hope) by some of the highly unusual features of Perdition’s Flames:

• You can’t die, because you’re already dead! The only time you’ll ever see “*** You have died ***” in Perdition’s Flames is at the very beginning.

• You can’t even get into an unrecoverable situation. This means that you can’t destroy important objects, you can’t do anything irreversible, and you can’t become trapped anywhere. We think that this sort of thing is a waste of your time, and doesn’t add anything to the enjoyment of the game—quite the opposite—so we left it out.

• You won’t need to eat or sleep. In our bold arrogance, we figured we could think of a puzzle or two that are more interesting than the eatingand-sleeping puzzle you’ve already seen in every other text adventure game.

• You won’t have to worry about how many items you’re carrying or how heavy they are. There might be puzzles involving heavy objects that you need help to lift, or places you can’t carry a particular object without solving a puzzle, but as a general rule, you won’t ever see an error message such as “You already have your hands full.”

• This game is certified 100% maze free!!! If you find yourself in a maze, please report it to High Energy Software at once and we will fix the error. If you really like mazes, we invite you to try your hand at another of our games, Deep Space Drifter.

Playing FLAMES

You control the game with commands in simple English. FLAMES will describe your surroundings, and you tell the game what you want to do.

Each time FLAMES shows you its prompt, >, you type a command—a simple imperative sentence. When you are done typing your command, press the RETURN (or ENTER) key; FLAMES doesn’t start interpreting your command until you press RETURN.

You can use capital or small letters, or even a mixture, in your commands; FLAMES doesn’t care. You can use words such as THE and AN when they’re appropriate, but you can omit them if you prefer.

You can abbreviate any word to six letters, but FLAMES will still pay attention to any more letters you type. For example, you could refer to a flashlight with the words FLASHL, FLASHLIG and so forth, but not with FLASHLASDF.

Travel

At all times during the game, you are in a particular location. FLAMES describes your location to you when you first enter, and again any time you type LOOK.

You can move from place to place in the story by typing the direction you want to go. FLAMES will always tell you the directions that you can go from your current location.

Generally, backtracking will take you back to where you started; for example, if you start off in a kitchen, go north to the living room, then go south, you will probably end up back in the kitchen. Occasionally, a path will be irreversible, but when this happens it’s usually made clear in the text why you can’t go back.

Most of the time, when FLAMES describes a door or doorway, you don’t need to open or close the door to go through the passage; the game will do this for you automatically. Only when FLAMES explicitly describes a closed door (or other impediment to travel) will you need to type a command to open it.

To move to a new location, type the direction you wish to go. FLAMES recognizes the eight compass directions (north, south, east, west, northeast, southeast, northwest, southwest), as well as up and down. You can abbreviate these to N, S, E, W, NE, SE, NW, SW, U, and D. In some locations you can also use IN and OUT.

Time

Time in FLAMES passes only in response to commands you type. Nothing happens while FLAMES is waiting for you to type something. Each turn takes about the same amount of time. If you want to let some time pass, because you think something is about to happen, you can type WAIT.

One thing to note about FLAMES is that there’s no particular time limit for completing the game. You’ll be dead for all of eternity, so there’s no hurry!

Score

FLAMES assigns you a score while you play, indicating how well you are doing. At certain points in the game, FLAMES will award you points when you accomplish something important. The score is intended to provide you with a measure of your progress in the game, and increases as you get further in the game; we’ve tried to assign points fairly uniformly through the game. You can’t lose points once they are earned.

The game will notify you whenever your score increases, by displaying a message like this:

      [Your score has just gone up.]

If you don’t want to see these messages, type NOTIFY and they will go away. If you want to bring them back, just type NOTIFY again.

Objects

In FLAMES’s imaginary world, you will find many objects that you can carry or otherwise manipulate. For example, you could type READ THE BOOK or OPEN THE DRAWER. Most of the objects in FLAMES have fairly obvious uses; you shouldn’t have to think of any obscure or unrelated words to manipulate an object in the story.

You usually won’t have to specify exactly where you want to put an object that you wish to carry; you can just type TAKE (followed by the object’s name) to carry an object.

Some basic verbs that you will often use are TAKE (to pick up an object), DROP (to drop an object), OPEN and CLOSE, and EXAMINE (which you can abbreviate to X), LOOK (which shows you the full description of your location; you can abbreviate LOOK to L), INVENTORY (which shows you the list of items you are carrying; you can abbreviate INVENTORY to I). In addition, you can sometimes PUT an object IN or ON another object.

Common Commands

We tried to make sure FLAMES uses only “obvious” verbs that naturally apply to the objects in the game: you push buttons, turn knobs, dig with shovels.

Here’s a list of some of the commands you can use in FLAMES. You probably won’t use any of these exact commands, since lots of the objects mentioned aren’t part of this game, but the list will give you an idea of how to phrase your commands. This list has an example of every verb and phrasing you will need to complete the game. You can use many other types of sentences as well, but you’ll never need anything that’s not on this list.

GO NORTH NORTH N UP TAKE THE BOX PUT THE FLOPPY DISK INTO THE BOX LOOK AT DISK CLOSE BOX TIE THE ROPE TO THE RAILING TYPE “HELLO” ON THE KEYBOARD TYPE 1234 ON THE KEYPAD WEAR THE HELMET TAKE OFF THE HELMET MOVE THE PILE OF LEAVES LOOK THROUGH THE PILE OF LEAVES SEARCH THE LEAVES OPEN CRATE WITH CROWBAR TURN ON THE LAMP TURN THE LAMP OFF STAND ON THE STOOL CLIMB UP THE LADDER CLIMB THE ROPE LIGHT MATCH TURN THE KNOB SIT ON THE CHAIR GET UP THROW BALL AT WINDOW MOVE THE LEVER UP TURN DIAL TO 15 GET IN THE CAR GET OUT OF THE CAR GET ON THE HORSE GIVE RAYGUN TO MARTIAN SHOW THE LETTER TO THE PROFESSOR ASK MARTIAN ABOUT PYRAMID ASK HAWKER FOR HOT DOG JUMP

Interacting with Characters

You’ll run into several characters during the game. You can interact with other characters in several ways. Some characters won’t be interested in talking with you at all, and they’ll usually let you know early on.

You can ask a character about something; for example, you could ASK MARTIAN ABOUT PYRAMID. You can also ask a character to give you something: ASK HAWKER FOR HOT DOG.

You can give a character an object (GIVE KEY TO JANITOR), or you can simply show something to a character (SHOW SWORD TO NINJA). These are often equivalent, but sometimes you can show someone an object to threaten or impress them.

Note that you can sometimes interact with characters in indirect ways. For example, if a character is standing on a trap door, and you pull the lever that opens the trap door, you’ll probably elicit some response from the character even though you didn’t use one of the usual character interaction commands.

Commands for Power Users

FLAMES understands much more complicated sentences than we’ve talked about so far.

You can usually use multiple objects in your sentences. Separate objects by the word AND or a comma:

TAKE THE BOX, THE FLOPPY DISK, AND THE ROPE PUT DISK AND ROPE IN BOX DROP BOX AND BALL

You can use the word ALL to refer to everything that is applicable to your command, and you can use EXCEPT (right after the word ALL) to exclude certain objects:

TAKE ALL PUT ALL EXCEPT THE DISK AND THE ROPE INTO THE BOX TAKE EVERYTHING IN THE BOX TAKE ALL ON SHELF

ALL refers to everything that makes sense for the command, except things inside containers. For example, if you are carrying a box and a rope, and the box contains a floppy disk, typing DROP ALL will drop only the box and the rope; the floppy disk will remain inside the box.

You can use IT and THEM to refer to the last object or objects that you used in a command, and HIM or HER to refer to the most recently mentioned characters:

TAKE THE BOX OPEN IT TAKE THE DISK AND THE ROPE PUT THEM IN THE BOX ASK MARTHA ABOUT THE MORTAGE GIVE HER THE SUBPOENA

You can put multiple commands on a single input line by separating the commands with periods or the word THEN, or with a comma or the word AND.

TAKE THE DISK AND PUT IT IN THE BOX TAKE BOX. OPEN IT. UNLOCK THE DOOR WITH THE KEY. OPEN IT,

AND THEN GO NORTH.

Each command will still count as a separate turn. If FLAMES doesn’t understand one of the commands on the input line, the game will tell you what it couldn’t understand, and it will ignore the rest of the commands on the line.

If you type a command that leaves out some important information, the game will try to figure out what you mean anyway. When only one object makes sense with the command, FLAMES will assume you meant that object:

>TIE THE ROPE (to the railing) The rope is now tied to the railing. The end of the rope nearly reaches the floor of the pit below.

If FLAMES can’t guess what you meant, the game will ask you for more information. You can answer by typing the missing information. If you decide you didn’t want to bother with the command after all, you can just type a new command; FLAMES will ignore the question it asked.

>UNLOCK THE DOOR What do you want to unlock the door with? >THE KEY Which key do you mean, the gold key, or the silver key? >GOLD The door is now unlocked.

Unknown Words

FLAMES will sometimes use words in its descriptions that it won’t recognize in your commands. For example, you may see a description such as, “The planet’s rings are are visible as a thin arc high overhead, glimmering in the sunlight.” If FLAMES doesn’t know words such as “sunlight,” you can assume that they’re not needed to progress through the game; they’re just described to make the story more interesting and vivid.

Saving and Restoring

During the game, you can save your position at any point to a disk file, then restore the same position later on. You can save as many different positions as you like.

To save the game, type SAVE at any command prompt; the game will ask you for the name of a disk file to use to store the game position. You must enter a filename suitable for your operating system, and the disk must have enough space to store the game state. FLAMES will tell you if the game could not be saved for some reason.

To restore a previously saved position, type RESTORE at the > prompt. The game will ask you for the name of a disk file; after the disk file has been read, the game state is restored to exactly the position it was when you saved it.

Special Features

FLAMES has several special features that can be helpful while playing. Refer to the reference card for the details of these features.

Command Line Editing: While you are typing a command, FLAMES allows you to edit the line you are typing. On most systems, you can simply use the left and right cursor-arrow keys to move the cursor to any point in the line. Then, insert new text at the cursor simply by typing the text. Use the BACKSPACE key to delete a character to the left of the cursor; most computers also have a DELETE key which allows you to delete the character after the cursor.

Command Recall: FLAMES remembers the several most recent commands you have typed, and allows you to recall them to enter again. At the command prompt, press the up cursor-arrow key to recall the previous command; press it again to recall the command before that, and so forth. Using the down cursor-arrow key reverses this process, until you get back to the original command you were typing before you started pressing the up cursor-arrow key. The game retains the several dozen most recent commands.

Once you have recalled a command, you can re-enter it by pressing the RETURN key. In addition, you can edit the command, as described above, before pressing RETURN.

Review Mode: At any time during play, you can go back and look at text that has scrolled off the screen by entering “review mode.” (On most systems, review mode is activated by pressing the function key F1; see your Installation and User’s Guide for the key on your computer.) The game will typically retain roughly twenty screens of text for review.

While in review mode, the normal status line is be replaced by a review mode help line, which shows you the keys you can use to view past text, and the key that exits review mode and resumes game play.

While you are in review mode, the screen becomes a window onto the text that FLAMES has stored away. When you first enter review mode, you are looking at the bottom of this text, which is the screenful of text that was just displayed. Use the up and down cursor-arrow keys to move the window up and down. Pressing the up cursor-arrow key moves the window up one line, showing you one line of text that has scrolled off the screen. Most systems also provide keys to move up and down by a full screenful (a “page”).

To exit review mode and resume game play, press the same key that you used to activate review mode.

Important Commands

FLAMES has several special commands that you can use to control the game. You can use these commands at any command prompt.

AGAIN—FLAMES will try to repeat your last command. If your last command line consisted of several sentences, only the last command on the line can be repeated this way. You can abbreviate AGAIN to G.

CREDITS—Displays a list of the people involved in writing the game.

INVENTORY—Shows the list of items you are carrying. You can abbreviate INVENTORY to I.

LOOK—Shows the full description of your location. You can abbreviate LOOK to L.

NOTIFY—Turns score notification on and off. At the start of the game, notification is on, which means that you’ll see a message each time your score increases. Just type NOTIFY to turn this feature off; if you decide you want it back after all, type NOTIFY again.

OOPS—Allows you to correct the spelling of a word in the last command. You can use OOPS when FLAMES displays this complaint in response to a command: “I don’t know the word “word.” Immediately after this message, you can type OOPS followed by the corrected spelling of the unknown word. You can only type one word after OOPS, so this command doesn’t allow you to correct certain types of errors, such as when you run two words together without a space. You can abbreviate OOPS to O.

QUIT—This stops the game. If you wish to save your position before quitting, you should use the SAVE command first.

RESTART—This starts the game over from the beginning.

RESTORE—Restores a position that you saved previously with the SAVE command.

SAVE—Stores the current state of the game in a disk file, so that you can restore your current position in the game later.

SCORE—Shows you your current score, the maximum possible score, and the number of turns you have taken so far.

SCRIPT—Starts writing everything you see on the screen (your commands and the game’s responses) to a disk file. FLAMES will ask you for the name of a file to use for the transcript. Use UNSCRIPT to stop making the transcript.

TERSE—Tells FLAMES that you wish to see only short descriptions of locations you have already seen as you enter them. The short descriptions generally consist of a few words describing the location, and a list of any items in the room (excluding immovable items). The game will give you the full description of a room only when you enter the room for the first time or explicitly type LOOK. The game starts off in TERSE mode. See also the VERBOSE command.

UNDO—Takes back the last command, restoring everything to the way it was before the command. You can use UNDO many times in a row to take back a series of commands; the maximum number of turns you can undo varies depending on your actual commands, but is typically about 100 turns.

UNSCRIPT—Turns off the transcript being made with the SCRIPT command.

VERBOSE—Tells FLAMES to show you the full description of every location as you enter, even if you’ve already seen the location. See also the TERSE command.

VERSION—Shows you the current version of FLAMES.

WAIT—Allows time to pass in the story without doing anything. When the game is waiting for a command, time does not pass in the game; you can use this command to wait for something to happen. You can abbreviate WAIT to Z.

Hint Book

If you get stuck during the game, you can turn to this section for help. We tried to think of all the questions you may have while playing the game, then answered each question with a decreasingly subtle series of hints. Each series of hints starts with a very slight nudge in the right direction—we tried not to give too much away in the early hints. The later hints become increasingly specific.

To use this section, look through the topic headings to find the general part of the game where you’re stuck. Then, look through the questions to see if you can find something relevant to your problem. When you find a suitable question, it will say something like “See hint 25.” That’s the first in the series of hints for the question. Turn to the page with hint 25, and read what it has to say. You’ll notice that most of the hints refer you to another hint—this is the next more specific hint in the series of hints for this question. Think about each hint, and if it doesn’t help, follow the reference to the next hint.

Some of the questions are irrelevant to the game. We’ve added these to make it less likely that the mere presence of a question gives away a puzzle— because some of the questions are “fake,” you shouldn’t assume that an object, place, or event mentioned in a question actually occurs in the game. Stick to the questions that seem to address the specific problems you’ve encountered.

We’ve divided the hints into two sections: regular hints, and “spoilers.” The spoiler section is separated out because these hints entirely give away puzzles by telling you exactly what to do. Unless you’re totally stuck on a problem, you’ll probably want to avoid reading the spoilers.

Questions

General Questions

What’s the point of the game? See Hint 372 Help! I keep dying at the very beginning of the game! See Hint 373 What did I die of? See Hint 303

The Boat Ride

How do I get off the boat? See Hint 131 How do I get a drink at the bar? See Hint 133 Infernolab

How can I get the dead bulb? See Hint 112 How can I win at the economics experiment? See Hint 370 How can I get something out of the dumpster? See Hint 193 How can I get into the dumpster? See Hint 38 How do I find a lighter? See Hint 12 What should I do with the flaming maintenance man? See Hint 361 How can I open the toolbox? See Hint 337 How do I get past the locked door in the lab? See Hint 357 What should I do at Infernolab? See Hint 2 How do I get rid of the technician? See Hint 107 What do I do once the technician is gone? See Hint 293 How can I get into the crawlways? See Hint 214 Why do I need to use the ID badge to exit the lab? See Hint 129

The Haunted House

What’s the portal? See Hint 308 Why do I keep going back to Hell when I leave the house area? See Hint 105 What do I do in the house? See Hint 156 I can’t do anything with any of the objects in the house! See Hint 299 The group of people keeps undoing what I do with the ring! See Hint 314 Okay, I’ve moved the ring! Now what? See Hint 182 The ring keeps falling off the porch! See Hint 7 Okay, the ring is in the styrofoam cup. What now? See Hint 27 How do I get the ring down the hall? See Hint 67 What do I do when the ring falls into the sink? See Hint 181 How do I get the ring out of the basement? See Hint 17 How do I get the ring away from the priest? See Hint 371

The Suburbs

How can I buy a mirror at the mirror store? See Hint 266 How can I buy an amulet at the amulet store? See Hint 136 What do I do with the dollar bill? See Hint 219 What can I do with the empty bag of Cheese Doodles? See Hint 176 How can I get the medicine cabinet open? See Hint 103 The mirror keeps breaking! See Hint 127 Is there anything interesting in the programmer’s office? See Hint 302 How can I get the ID card from the programmer? See Hint 221 I can get the ID card, but the programmer keeps taking it back from me! See Hint 75 What can I do with the programmer? See Hint 215 The Mountain Lake

Is there anything I can do with the water? See Hint 319 How can I get into the underwater cave? See Hint 134

The Lava River

I keep missing the stalagmite! See Hint 212 How can I land on the shore along the lava river? See Hint 146 What can I do with the stalagmite? See Hint 244

Heaven

How can I go to Heaven? See Hint 21 How can I cash the check? See Hint 367 How do I get past the waiting room? See Hint 220 I need more money to pay for a meal. See Hint 316

The Desolation Area

What can I do at the bottom of the crater? See Hint 83 How do I get past the lake of blood? See Hint 30 Where do I find the bracelet? See Hint 91

The Bridge

How do I get past the bridge? See Hint 249

The Department of Motor Vehicles

How can I get a driver’s license? See Hint 324 Where do I find the appointment card? See Hint 343 How do I get a pencil? See Hint 277 How can I sharpen the pencil? See Hint 87 What are the speed limits on the driver’s license test? See Hint 69 How do I solve the first math problem on the driver’s test? See Hint 188 How can I solve the second math problem on the driver’s test? See Hint 47

The Celestial Security Office

How can I get rid of the other applicant? See Hint 321 How can I get the applications out of the waste basket? See Hint 235 How can I fill out the applications? See Hint 36 How can I get my Celestial Security Card? See Hint 25 How can I get past the employee and into the room to the west? See Hint 276 What do I do with the machine? See Hint 5

The Island

How do I get past the thicket on the island? See Hint 54 How can I get into the pyramid? See Hint 144 Is there some kind of trap if I take the gold skull? See Hint 246 What can I do about the pterodactyl? See Hint 247

Hints

1. Have you looked at the gap in the statues after moving the lion-headed statue? Next hint: 448 (Spoiler)

2. There’s an economics experiment going on. Next hint: 294

3. If he didn’t see you pick it up, he wouldn’t take it back. Next hint: 466 (Spoiler)

4. The amulet store is there for you to learn about the amulets.

5. Have you read the display? Next hint: 200

6. You could wait until the fire goes out. Next hint: 65

7. Think about how you got it out there in the first place. Next hint: 169 8. A portion of the sphere sticks outside the pyramid. Each face of the pyramid cuts through part of the sphere. Next hint: 48

9. As a result of being able to manipulate the group, you can manipulate certain objects: every time you go somewhere that involves a door, the group of people will have to open the door.

10. Have you looked at it carefully? Next hint: 360

11. So long, in fact, that even eternal beings might die and start rotting? Next hint: 336

12. Have you run into anyone that might have one? Next hint: 108

13. Is there anything small and mobile that you’ve encountered? Next hint: 410 (Spoiler)

14. She wants to get a driver’s license. Next hint: 46

15. A parachute would help. Next hint: 55

16. Have you been in the alley next to Infernolab? Next hint: 283

17. This one’s easy. Next hint: 230

18. Now you need the combination to the locker. Next hint: 259

19. We wish to integrate the volume of the sphere outside the pyramid. We must therefore determine the minimum distance from a face of the pyramid to the center of the sphere, and integrate from that distance to \frac{a}{2}. Next hint: 26

20. Have you found anything useful for prying things open? Next hint: 338 21. Have you noticed the Space Needle? Next hint: 39

22. Perhaps the crowbar would help. Next hint: 284

23. So long, in fact, that even eternity might start seeming short? Next hint: 11

24. The Celestial Security Administration recycles its paper. Next hint: 93

25. If you could get rid of the other applicant, you would be next in line. Next hint: 250

26. The triangle formed by the height of the pyramid, a line from the center of the sphere to the center of a base edge, and the line along the face connecting them will have two known sides: the height, \frac{a}{\sqrt{2}}, and the line from the center to the base edge: \frac{a}{2}. The length of the line down the center of the face therefore has length a\frac{\sqrt{3}}{2}. Next hint: 40

27. Did you fill the cup with coffee? Next hint: 138

28. Actually, you can’t reach it, but you might have something that can. Next hint: 60

29. You need to get past the locked door in the Infernolab hallway first. Next hint: 18

30. Have you heard of anything that might protect you? Next hint: 41 31. You’re not strong enough to swim against the current without some kind of assistance. Next hint: 335

32. The robot is in the insurance office. Next hint: 317

33. What does he ask for when you try to take the paper while carrying the bin? Next hint: 208

34. It would be great if there were a mirror somewhere above the partitions. Next hint: 152

35. Have you found anything useful for prying things open? Next hint: 449 (Spoiler)

36. They’re all crumpled up, making them difficult to fill out. Next hint: 353

37. You can’t do anything with the dollar bill. They only accept the local currency, “hellmids,” in Hell.

38. Have you tried climbing in? Next hint: 190 39. There’s an elevator in the Space Needle. Go look at its controls. Next hint: 325

40. The shortest line from the center to a face will be perpendicular to the face. We can exploit our knowledge that the faces are equilateral triangles to determine that the normal from the center of the sphere will intersect the face at its center. We can thus determine that the line from the center of a base edge to the center of the face is of length \frac{a}{2\sqrt{3}}. From this, and the knowledge that the distance from the center of the sphere to the center of a base edge is \frac{a}{2}, we can deduce that the distance from the center of the sphere to a face is \frac{a}{\sqrt{6}}. Next hint: 270

41. Have you been to the amulet shop? Next hint: 281

42. Have you found the access panel? Next hint: 143

43. Have you found anything useful for extinguishing fires? Next hint: 238

44. Look at the slots. Next hint: 352

45. If you haven’t already been into the crawlways, don’t go on until you have. Next hint: 186

46. What did you need to get a driver’s license? Next hint: 333

47. I’ll bet you thought we were going to leave this as an exercise to the reader! But not so: this is a full-service hint book, prepared with the latest mathematical typesetting software. Square roots and integrals don’t scare us one bit! Next hint: 8

48. This problem requires calculus. Next hint: 354

49. Have you noticed how to lose the group? Next hint: 51

50. Have you tried standing on the table? Next hint: 358

51. If you go from room to room, they’ll follow you. Next hint: 417 (Spoiler)

52. A moving target would be best. Next hint: 13

53. Unfortunately, he carries it everywhere he goes. Next hint: 350

54. You’ll need something to help you break through it. Next hint: 192 55. Where might you find a parachute? Next hint: 275

56. A line drawn from the center of the sphere to the point at which an edge of the pyramid intersects the sphere is a radius, so it has length \frac{a}{2}. Consider the triangle formed by this radius, the portion of the edge from the intersection point down to the base, and a line from there to the center of the sphere. We can deduce that the third line (from the center of the sphere to a corner of the base) has length \frac{a}{\sqrt{2}}. Therefore, the second line (the line along the edge) has length \frac{a}{2}. Next hint: 413 (Spoiler)

57. What’s more, she’ll never finish. Next hint: 96

58. Have you found the radio? Next hint: 278

59. You need to put out the fire. Next hint: 43

60. How about the rope? Next hint: 128

61. She’s actually looking for something she lost. Next hint: 359

62. However, you still need to fool the official into thinking that you’re going to recycle the paper. Next hint: 33

63. Did you turn on the faucet? Next hint: 310

64. You’ll need a sharp pencil (that’s another question). Next hint: 139 65. But that would be pointless, since it never will. Next hint: 59

66. Perhaps you could extend your flight somehow. Next hint: 229

67. This is similar to what you had to do with the trap door. Next hint: 201

68. You can put things in the bag of cheese doodles. Next hint: 113

69. You can hear about speed limits somewhere. Next hint: 58

70. Maybe you could force it open with something. Next hint: 168

71. The main way to interact with him is to break the mirror. Next hint: 119

72. But you’ll need to practice throwing the rope to be any good. Next hint: 272

73. You can’t climb the building. Next hint: 45

74. Have you seen any alcohol anywhere? Next hint: 260

75. Have you noticed why he keeps taking it back? Next hint: 202

76. Maybe you could ask the owner about it. Next hint: 124

77. You’ll need to find another way to get your Celestial Security card. Next hint: 348

78. You need a driver’s license and a Celestial Security Card.

79. You could try to take it from him. Next hint: 329

80. Have you looked at the pipe valve at the west end of the crawlway? Next hint: 109

81. If you try it, you’ll find there’s not a trap after all.

82. You should also notice that you have a certain ability to manipulate the group of people. Next hint: 320

83. Have you noticed the big stone slab? Next hint: 344

84. You can’t buy anything at the mirror store.

85. Maybe the battery is dead. Next hint: 111

86. There’s something already missing in the cockpit. Next hint: 479 (Spoiler)

87. There’s a pencil sharpener in the programmer’s office. Next hint: 327 88. What have you found that might fit the holes? Next hint: 467 (Spoiler)

89. Something on the island. Next hint: 364

90. You can’t distract the teenager. Next hint: 84

91. It’s someplace hard to get to. Next hint: 211

92. The equipment to make cards is in the back office. Next hint: 158 93. Have you seen any recycling equipment anywhere? Next hint: 346 94. The volume of the whole sphere is \frac{4}{3}π(\frac{a}{2})^{3}, so the volume of a hemisphere is half that: \frac{4}{3}π \frac{a^{3}}{16} , or π \frac{a^{3}}{12} . Next hint: 291

95. You need to find an alternative entrance. Next hint: 178

96. You need to find another way to get a card. Next hint: 205

97. Did you look in the styrofoam cup? Next hint: 173

98. You can get beer or liquor: ASK BARTENDER FOR A BEER, for example.

99. Have you gone flying through the air yet? Next hint: 154

100. The programmer has the ID badge. Next hint: 79

101. You’ll have to find another way to get your Celestial Security Card.

102. Do what he did. Next hint: 456 (Spoiler)

103. It’s stuck. You need something to improve your leverage. Next hint: 20 104. And there’s no way to get it in there to begin with!

105. Maybe there’s some scientific reason that you can only exist as a ghost within the immediate vicinity of the dimensional gateway. Next hint: 267 106. Have you tried walking through a wall? Next hint: 184

107. Did you ask the programmer about Infernolab? Next hint: 177

108. The sullen teen in the mirror store seems to smell of cigarette smoke. Next hint: 117

109. Open the valve, wait a few turns, then light a fire. Next hint: 137 110. The dollar bill isn’t of any use. Next hint: 254

111. Maybe it would work if you replaced the battery with a working one. Next hint: 197

112. Did you notice who put it there? Next hint: 262

113. That’s all it’s good for (in other words, it’s not particularly useful).

114. Maybe you could find a way to see other cubicles using the access panel. Next hint: 34

115. The machine makes the card you’re looking for. Next hint: 216

116. Have you followed him to the mirror store? Next hint: 457 (Spoiler)

117. And, if you watch carefully, you might notice he seems to want a cigarette. Next hint: 171

118. Have you found the ID badge? Next hint: 100

119. When you break the mirror, the programmer goes and buys a new mirror to replace the broken one. Next hint: 483 (Spoiler)

120. You can swim in it.

121. Ask the bartender for a drink. Next hint: 98

122. You can only get it when he’s not carrying it. Next hint: 53

123. Have you noticed what happens when you break the mirror? Next hint: 126

124. Of course, as long as he’s on fire, he won’t talk to you. Next hint: 6 125. He’s guarding a couple of items you’ll need. Next hint: 362

126. Have you followed the programmer to the mirror store? Next hint: 196 127. Have you noticed why it breaks? Next hint: 207

128. Throw the rope at the stalagmite. Next hint: 72

129. The building codes in Hell are not very well enforced. Next hint: 165 130. Have you found your way into the crawlway? Next hint: 164

131. Wait until it docks Next hint: 217

132. The pterodactyl evidently is very protective of the gold skull. Next hint: 222

133. Look at the bar for a selection of drinks. Next hint: 121

134. There’s a strong current dragging you away. Next hint: 31

135. Stand on the table. Look around. Look at the railing. Next hint: 486 (Spoiler)

136. Have you asked for a free sample? Next hint: 297

137. You need the lighter (see the questions about that). Next hint: 455 (Spoiler)

138. Probably not, because there is no coffee to fill it with. Next hint: 104

139. You’ll also need to figure out about speed limits so you can pass the test (see the question about speed limits for help there).

140. Have you noticed anything that might explode in the crawlway? Next hint: 80

141. Have you seen any electronic items that might contain batteries? Next hint: 340

142. Besides, this was a fake question to begin with!

143. Open it and then go to the cubicle and look around. Next hint: 306 144. Have you tried entering it? Next hint: 298

145. One is the crowbar. Next hint: 179

146. You can’t when you’re swimming in the river. Next hint: 305

147. Did you look carefully around the barn area? Next hint: 273

148. Since Hospital Zones and 43 mph zones are mentioned in one sentence as different zones, you can assume that Hospital Zones are not 43 mph. Next hint: 257 149. Even if you could fill them out, you can’t get rid of the other applicant, so you’d never have a chance to submit your application. Next hint: 77 150. The maintenance worker at the lab has a uniform. Next hint: 356 151. But you can make it better. Next hint: 334

152. Maybe you could put something over the access panel. Next hint: 463 (Spoiler)

153. Don’t read the next hint until you’ve obtained the flippers. Next hint: 409 (Spoiler)

154. If so, have you noticed where you land? Next hint: 331

155. Have you talked to God? Next hint: 447 (Spoiler)

156. You’re a ghost! Have some fun haunting these people! Next hint: 106

157. The compartment is stuck. You need something to force it open. Next hint: 35

158. However, you can’t get past the officer. You must find another entrance. Next hint: 99

159. Have you seen anything that might be able to withstand the heat? Next hint: 282

160. You can’t reach anything in the dumspter from outside the dumpster. Next hint: 301

161. Have you found a way to distract the programmer? Next hint: 123 162. The technician also mentioned alcohol a couple of times. Next hint: 74

163. Try moving them around. You need to move both of them to make something happen. You can move them both up, both down, or one up and one down. Next hint: 404 (Spoiler)

164. Have you noticed any explosions? Next hint: 140

165. It’s designed that way for the same reason that some cars make you use your key to lock the door from the outside: you can’t lock your ID badge behind the door if you need it to get out. It’s supposed to be impossible to get stuck in this game, remember.

166. You need the application forms from the waste basket to figure out what the questions on the display mean.

167. Actually, you need a single object to put in the holes. Next hint: 315 168. Have you found anything useful for prying things open? Next hint: 22

169. Did you find the car on the driveway? Next hint: 243

170. There’s a bucket for transporting water. Next hint: 419 (Spoiler)

171. If you offer him a cigarette, maybe he’ll be helpful. Next hint: 279 172. Some object might be helpful in extending your flight. Next hint: 15 173. What are you talking about? You can’t get it out there in the first place!

174. Have you tried landing on the east side of the volcano? Next hint: 236

175. In fact, there’s no way to go from the Celestial Security Office front room to the back room. Next hint: 194

176. Don’t go on until you have the airplane working. Next hint: 318 177. He mentioned that the workers at the lab are a bunch of drunkards. Next hint: 162

178. The alternative entrance isn’t from the ground. Next hint: 289

179. The other is the fireplace poker. Next hint: 163

180. Hence, the area of a cross-sectional circle at radius r out from the center will be π(\frac{a^{2}}{4} − r^{2}). We therefore must integrate ∫ π(\frac{a^{2}}{4} − r^{2})dr. Next hint: 19

181. Did you close the sink trap? Next hint: 63

182. Did you notice where it fell? Next hint: 195

183. Have you found any ways to make the programmer leave for a while? Next hint: 418 (Spoiler)

184. Or just disappearing into the ground? Next hint: 292

185. You can put it in the bucket. Next hint: 120

186. Don’t go on until you’ve figured out how to go from the crawlways directly to the courtyard. Next hint: 347

187. There’s a drawer in the desk. Next hint: 224

188. Since the sides of the base are tangent to the sphere, and the base is a square, and the center of the sphere is in the plane of the base, the sides of the base must be the same length as the diameter of the sphere. Next hint: 256 189. Have you been in the back office? Next hint: 92

190. The dumpster is too tall. You can’t enter it from the alley. Next hint: 300

191. Maybe you should put things in the holes. Next hint: 167

192. Something big. Next hint: 89

193. Have you tried just taking what you want? Next hint: 160

194. However, there’s a different way to get there. Next hint: 95

195. It fell on the trap door. Next hint: 241

196. Have you noticed what happens when the programmer pays for the mirror? Next hint: 443 (Spoiler)

197. Have you seen any other batteries? Next hint: 141

198. You can get twenty hellmids by winning at the economics experiment. Next hint: 341

199. You need to give something to the troll. Next hint: 263

200. You can type responses on the keyboard with commands such as TYPE

"MIKE" ON KEYBOARD. Next hint: 115

201. However, you’ll notice that the group of people will notice the ring (and return it to its original place on the trunk) if you lead them out the front door while the ring is in front of it. Next hint: 475 (Spoiler)

202. It’s because he sees you pick it up. Next hint: 3

203. There’s always UNDO. Next hint: 304

204. Since both School Zones and 27 mph zones are mentioned in one sentence, they’re not the same. Likewise for Library Zones and 43 mph zones. In addition, because of what you learned about sign sizes, you can eliminate School Zones from being 43 mph and Library Zones from being 27 mph. Next hint: 223

205. Maybe you could make your own card. Next hint: 189

206. Now, imagine you were doing this hobby or recreational activity, and something went terribly wrong, and you died.

207. It breaks because it falls onto the sink. Next hint: 309

208. You need a uniform. Next hint: 312

209. Did you notice any places where things might go missing in the cockpit? Next hint: 86 210. Did you listen to what the lecturer in the immigration center said to the lobbyist? Next hint: 242

211. It’s hidden. Next hint: 153

212. Your aim isn’t very good. Next hint: 151

213. To get a ticket, you need twenty hellmids. Next hint: 198

214. Did you notice what the maintenance man did in the lobby? Next hint: 351

215. You can ask him about a few things. Next hint: 125

216. You simply need to enter the information requested by the machine. Any set of information will do, as long as it’s internally consistent. Next hint: 166

217. It won’t dock until after you’ve been into the control room.

218. Maybe it’s blocked by something. Next hint: 438 (Spoiler)

219. You could buy insurance from the robot. Next hint: 32

220. Have you talked to the woman? Next hint: 286

221. You can’t get it from him directly. Next hint: 122

222. Maybe you should find something that it would want to protect even more. Next hint: 1

223. We now learn that Residential Zones are either 17 or 23 mph. This means that Residental Zones are not 43, 59, or 27 mph. Next hint: 288 224. Have you tried opening the drawer? Next hint: 355

225. Maybe there’s one in the programmer’s office somewhere. Next hint: 228

226. Keep in mind that there are only five different types of zones, and each zone has a different speed limit. Next hint: 148

227. You’ll need to go explore the wastelands in western hell. Next hint: 489 (Spoiler)

228. Well, there isn’t one just lying around, but maybe it’s somewhere other than out in the open. Next hint: 187

229. There’s an object that you can use to extend your flight. Next hint: 440 (Spoiler)

230. Almost anything you do to the ring will get it out of the basement. Next hint: 414 (Spoiler)

231. Perhaps the owner has the key. Next hint: 76

232. There’s a lake nearby. Next hint: 170

233. Have you been to the suburban house? Next hint: 225

234. Of course, there is no hole in the rock, and you can’t land on the east side of the volcano, and you can’t get the airplane working. Next hint: 68 235. Notice what happens when you try to take them. Next hint: 24 236. Did you look in the hole in the rock? Next hint: 234

237. You can give her the one you used. Next hint: 255

238. Have you noticed water anywhere? Next hint: 326

239. You should find something to practice on. Next hint: 52

240. You need to find one somewhere else. Next hint: 258

241. Maybe if you could open the trap door, the ring would fall through. Next hint: 311

242. You still need to determine which speed limit belongs to which zone. You can do this using a simple deductive process. Listen to the various speed limits and zones mentioned, and rule out each impossible combination. Using the information presented, you’ll be able to rule out all but one possible combination of zones and speeds. The easiest way to do this is to make a grid, with zone names for columns and speed limits for rows. Next hint: 226

243. The keys for the car are somewhere in the house. Next hint: 97

244. Maybe you could reach it. Next hint: 28

245. We learn that either School Zones or Hospital Zones are are 23 mph. We already know that Hospital Zones are 27 mph, so it must be School Zones that are 23 mph. This tells us that Residental Zones are 17 mph. We now know all the speed limits.

246. Why don’t you try it? Next hint: 203

247. Didn’t this sound like a fake question? Next hint: 132

248. You need objects that are long and narrow. Next hint: 145

249. Read the junk mail; it contains a clue about this. Next hint: 199 250. Unfortunately, you can’t get rid of the applicant until she’s done. Next hint: 57

251. Have you noticed the compartment? Next hint: 157 252. That won’t work, unfortunately. Next hint: 142

253. But it’s dry. At least you might find the bucket useful. Next hint: 313

254. In fact, there’s no more money to be had at all. Next hint: 265

255. She’ll appreciate it a lot; however, that’s not why she’s here. Next hint: 61

256. Therefore, the sphere has radius \frac{a}{2}. Next hint: 56

257. Since signs in School Zones are blue and signs in 59 mph zones are green, you know that School Zones are not 59 mph zones. Next hint: 204 258. Does it seem like the DMV keeps people waiting a long, long time? Next hint: 23

259. What did the maintenance worker have when you first saw him? Next hint: 471 (Spoiler)

260. How about in the bar on the boat? Next hint: 452 (Spoiler)

261. Try climbing the spikes. Next hint: 44

262. Did you notice how he put it there? Next hint: 102

263. You need a great treasure. Next hint: 271

264. You need the key to the toolbox. Next hint: 231

265. You need to find an alternative form of payment. Next hint: 342 266. Have you tried distracting the teenager? Next hint: 90

267. But the actual scientific reason is that we didn’t feel like implementing anything more in this section, and it seemed like an easy way to keep you from straying too far.

268. Get in the booth.

269. Something small that’s precariously balanced. Next hint: 82

270. We integrate ∫ \frac{a}{2} ; \frac{a}{\sqrt{6}} π(\frac{a^{2}}{4} − r^{2})dr and find we must evaluate π(\frac{ra^{2}}{4} − \frac{r^{3}}{3} ) from \frac{a}{\sqrt{6}} to \frac{a}{2}. This value is πa^{3}(\frac{1}{4} − \frac{1}{12\sqrt{6}}). Next hint: 453 (Spoiler)

271. The treasure is from the oldest part of Hell. Next hint: 458 (Spoiler)

272. You need to find another small, moving object to practice with. Next hint: 436 (Spoiler)

273. How about in the cockpit? Next hint: 209 274. Have you found your way into the crawlway? Next hint: 42

275. The airplane might be a good place to look. Next hint: 251

276. You can’t distract or otherwise get rid of the employee. Next hint: 290

277. Pencils are usually around offices. Next hint: 233

278. You need to listen to the public affairs station on the radio. Next hint: 210

279. Have you seen any signs in the desolation area to the west? Next hint: 487 (Spoiler)

280. Actually, though, you might notice that you can “nudge” objects when you try to manipulate them. Next hint: 330

281. You need amulets against various threats. Next hint: 323

282. Have you been past the locked door in the lab? Next hint: 454 (Spoiler)

283. Have you looked at the dumpster? Next hint: 287

284. Unfortunately, the crowbar is inside the toolbox. Next hint: 264 285. Have you tried looking over the partition? Next hint: 50

286. Do you know why she’s come here? Next hint: 14

287. There’s a pair of flippers in the dumpster. Next hint: 403 (Spoiler)

288. We can now see that Freeways must be 43 mph, since no other zones are left to be 43 mph. This means that no other zone is 43 mph, which means that Hospital Zones must be 27 mph. This tells us that Library Zones must be 59 mph. Next hint: 245

289. You need to get onto the Celestial Security office roof. Next hint: 73 290. You can’t fool him into thinking you’re authorized. Next hint: 175 291. To integrate the volume of a sphere, integrate the area of a crosssection along a radial line. A cross-section at distance r out from the center of the sphere has radius \sqrt{\frac{a^{2}}{4} − r^{2}}. (We can deduce this from drawing the right triangle formed by a radius of the sphere, our distance out from the center of the sphere, and a line normal to this second line out to the sphere. The length of the first line is r, the second is \frac{a}{2}, so the third—the item of interest—is easily derived.) Next hint: 180

292. Or going outside? Next hint: 412 (Spoiler)

293. There’s a button and a booth. Next hint: 368

294. There’s also some interesting stuff beyond the door with the card reader. Next hint: 366

295. Do what they did. Next hint: 407 (Spoiler)

296. But the programmer won’t let you as long as he’s there. Next hint: 328

297. There’s no way to ask for a free sample. Next hint: 369

298. Maybe there’s an entrance around here somewhere. Next hint: 218 299. Of course not! You’re a ghost, for crying out loud! Next hint: 280 300. You need to enter it from above somehow. Next hint: 130

301. You need to get into the dumpster (see the question about that for more hints).

302. There’s a drawer that you might want to look inside. Next hint: 296 303. Pick your favorite dangerous hobby or recreational activity. Next hint: 206

304. Besides, nothing in this game can kill you. Next hint: 81

305. You’ll need something to use as a boat. Next hint: 159

306. The access panel is directly over the cubicle. Next hint: 114

307. Have you noticed anything that distracts the programmer? Next hint: 472 (Spoiler)

308. It’s a dimensional gateway. If you enter the portal, you will return to Hell. But you can come back as often as you want.

309. Maybe you could cushion its fall somehow. Next hint: 476 (Spoiler)

310. I don’t believe it ever got into the sink!

311. Of course, you can’t open the trap door directly, but you might be able to get someone else to do it. Next hint: 430 (Spoiler)

312. A maintenance worker uniform will do. Next hint: 150

313. There’s a lake in the mountains. Next hint: 469 (Spoiler)

314. That’s because they think it’s terribly interesting what you’re doing. Next hint: 363

315. What sort of object would persons in authority in Hell carry? Next hint: 88 316. You can’t participate in the economics experiment again. Next hint: 110

317. The insurance office is in Ditch Day Drifter, another fine TADS game available from High Energy Software. Next hint: 37

318. Have you noticed anything flying over the volcano? Next hint: 174 319. You can drink it. Next hint: 185

320. At least, you can make them go somewhere by going there yourself. Next hint: 9

321. You can’t. Next hint: 101

322. The answer is in some junk mail you’ll find in a suburban home. Next hint: 478 (Spoiler)

323. The amulets are found all around Hell in hard-to-reach places. Go to the amulet shop to find out which amulet is which. You’ll need the original of each of the amulets in the shop.

324. You need an appointment card (see the question about that for more hints). Next hint: 64

325. You need a ticket. Next hint: 213

326. There’s a well on the farm. Next hint: 253

327. But it doesn’t seem to work. Next hint: 10

328. You need to get the programmer out of his office somehow. Next hint: 307

329. He won’t give it to you, though, so you’ll have to find another way. Next hint: 161

330. There might be some object that would be moved by a nudge. Next hint: 269

331. Perhaps you could extend your flight somewhat. Next hint: 172 332. Have you tried saying “Get out”? Next hint: 252

333. She needs an appointment card. Next hint: 237

334. All you need is some practice. Next hint: 239

335. Maybe you could find something that helps you swim better. Next hint: 16

336. Have you seen any bones lying around? Next hint: 227 337. It’s locked. Next hint: 70

338. Have you found the crowbar? Next hint: 408 (Spoiler)

339. Water usually puts out fires. Next hint: 232

340. Anything portable? Next hint: 474 (Spoiler)

341. You can buy a ticket with twenty hellmids (cash only) at the vending machine in the Civic Center courtyard.

342. A valuable discount coupon could be helpful here. Next hint: 155 343. You can’t get it from the DMV. Next hint: 240

344. Did you see the three holes? Next hint: 191

345. Have you noticed anything he’s carrying? Next hint: 71

346. How about the blue bin in the suburban kitchen? Next hint: 62 347. Have you noticed that you fall just short of the Celestial Security building when you land in the courtyard? Next hint: 66

348. It involves getting into the back room. See the hints about that.

349. Have you noticed the lion-headed statue? Next hint: 261

350. However, he might accidentally lose it momentarily. Next hint: 116 351. Do what he did and pay careful attention. Next hint: 135

352. You need something to stick in the slots. Next hint: 248

353. You can’t change that. In fact, you can’t fill out the applications. Next hint: 149

354. We must integrate the volume of the portion of the sphere outside the pyramid, then subtract that from the volume of a hemisphere. Next hint: 94

355. The programmer won’t let you while he’s there. Next hint: 183

356. He’ll leave it in a locker when he’s done with it. Next hint: 29

357. Have you noticed the slot next to the door? Next hint: 118

358. You can’t look over the partitions, so you’ll have to find another way. Next hint: 274

359. Maybe you could find what she lost. Next hint: 147

360. Have you noticed the compartment? Next hint: 365

361. Maybe you should extinguish the fire. Next hint: 339 362. You might want to get into his desk. Next hint: 345

363. But if they weren’t there, you could do as you wanted. Next hint: 49 364. Have you looked at the statues? Next hint: 349

365. Have you looked in the compartment? Next hint: 85

366. Pay attention to what was going on with the maintenance workers. Next hint: 295

367. You need two forms of ID. Next hint: 78

368. Press the button. Next hint: 268

369. You can’t buy anything at the amulet store. Next hint: 4

370. It sure would be nice if you could see what the other subjects were doing before you made your selection. Next hint: 285

371. You have to scare him somehow. Next hint: 332

372. You’ll figure that out eventually. The answer is in the suburbs. Next hint: 322

373. The game takes place in Hell. How else do you suppose you’re supposed to get there other than by dying?

Spoilers

401. Just give or show the egg to the pterodactyl when it swoops down on you. (You have just one turn to do this; you must do it immediately after picking up the skull.)

402. Wear the gloves, then move the animal bones. Look at the skeleton.

403. You need to get the flippers and wear them, then go swimming.

404. The crowbar and poker will act as controls to steer the feet. The right lever steers the right foot, the left lever the left foot. Pushing a lever up steers the foot forward, down steers back. Next hint: 464 (Spoiler)

405. How about rubber gloves? Next hint: 446 (Spoiler)

406. Find the group, and lead them to the attic. They’ll open the trap door, and the ring will fall to the top of the stairs.

407. Stand on the table. Have a look around. Look at the railing. Next hint: 480 (Spoiler)

408. Pry the medicine cabinet open with the crowbar. Next hint: 433 (Spoiler)

409. It’s in the same place as the flippers. Next hint: 416 (Spoiler)

410. How about in the field? Next hint: 465 (Spoiler)

411. You need to go into the “Game Intervention” menu and select “Today’s Special.” This will yield a valuable discount coupon that you can use to pay for the meal.

412. Did you go up to the attic? There’s something interesting there.

413. The radius line is perpendicular to the edge, because the edge is tangent to the sphere at that point. Therefore, we have a right triangle with equal legs, meaning that the leg angles are 45 degrees. Hence, the triangle formed by our radius line, the other half of the edge, and the line from the center of the sphere to the top of the pyramid, is a right triangle with equal legs. We already know that one of the legs (the radius) has length \frac{a}{2}, hence the other leg (the edge) has length \frac{a}{2}. Hence, the height of the pyramid is \frac{a}{\sqrt{2}}, or a\frac{\sqrt{2}}{2}, as some people prefer to write it.

414. For example, just try to pick it up. Next hint: 427 (Spoiler)

415. Just go outside, then reenter the house somewhere else, and avoid the group of people. Then go up to the attic and deal with the ring.

416. Did you look at the trash? Next hint: 470 (Spoiler)

417. However, if you go outside, they won’t know where you went. They’ll never follow you into the meadow. They’ll also never follow you underground, obviously. Next hint: 415 (Spoiler)

418. How about breaking the mirror? Next hint: 422 (Spoiler)

419. Get the bucket, go fill it with water, then throw the water on the man.

420. However, if you wait for him to leave, you’ll be able to take it.

421. The portable radio in the farmhouse has a good battery. Put its battery in the sharpener, and the sharpener will function.

422. While he’s gone, you can look through his things.

423. The crowbar is in the toolbox. Next hint: 461 (Spoiler)

424. There’s something interesting underneath, if you could just get to it. Next hint: 432 (Spoiler)

425. Now you need to rotate the statue so it will walk into the thicket. Push the left lever up and the right lever down. The statue will rotate to face north. Next hint: 450 (Spoiler)

426. Look at the hole. Next hint: 485 (Spoiler)

427. When you try to do something with the ring, the priest will do exactly what you want. Next hint: 441 (Spoiler)

428. You can also follow him to the mirror store. Something interesting might happen there.

429. It’s the pitchfork. Just put the pitchfork in the holes.

430. The group of people will open the trap door if they know you’re in the attic. Next hint: 406 (Spoiler)

431. Try looking through the trash.

432. You need something to protect you against electrical shocks. Next hint: 445 (Spoiler)

433. The crowbar is in the toolbox.

434. See the questions about the stalagmite.

435. You can get a drink from the bar on the boat (just ask the bartender for a drink).

436. Have you run into any small animals that run around a lot? Next hint: 481 (Spoiler)

437. It’s something you will find in the barn. Next hint: 429 (Spoiler)

438. Something that you put there in the first place, in fact. Next hint: 444 (Spoiler)

439. Practice with the rabbit. Once you’re good enough, you’ll be able to hook the rope around the stalagmite.

440. The parachute will do the trick. Next hint: 488 (Spoiler)

441. He’s doing it to get rid of you, though. It would be courteous to oblige him and leave. Next hint: 477 (Spoiler)

442. Have you noticed the heat shield? Next hint: 451 (Spoiler)

443. Follow the programmer to the mirror store, wait until he drops his ID badge and leaves, then pick up the badge.

444. It’s blocked by that statue standing in front of the pyramid! Next hint: 468 (Spoiler)

445. Insulation of some kind might be useful. Next hint: 405 (Spoiler)

446. There are some rubber gloves in the suburban house’s kitchen. Next hint: 402 (Spoiler)

447. Have you explored the whole menu tree? Next hint: 411 (Spoiler)

448. Go through the gap, and climb the wall. You’ll find something useful. Next hint: 401 (Spoiler)

449. The crowbar will help here. Next hint: 423 (Spoiler)

450. Now push both levers up, and the statue will walk through the thicket.

451. Get the heat shield. Put it in the lava, then get in it. Next hint: 434 (Spoiler)

452. Give the technician a drink. Next hint: 435 (Spoiler)

453. We now subtract 4 times this quantity from the volume of a hemisphere and find the result: πa^{3}(\frac{7\sqrt{6}}{54} − \frac{1}{4}).

454. Have you been into the missile lab? Next hint: 442 (Spoiler)

455. Set off an explosion, then return to the crawlway. Next hint: 426 (Spoiler)

456. Stand on the table.

457. Follow him to the mirror store, and wait for him to pay for the mirror. He’ll drop it when he’s getting money out of his pocket. Next hint: 460 (Spoiler)

458. It’s from the island past the lake of blood. Next hint: 484 (Spoiler)

459. Now you just need to know what the line numbers on the application form mean. For this, you’ll need an application form.

460. If you pick it up while he’s in the room, he’ll notice and snatch it back from you. Next hint: 420 (Spoiler)

461. Wear the parachute and set off the explosion. Pull the red rip-cord once you are airborne. You’ll land on the roof of the Celestial Security office. Next hint: 459 (Spoiler)

462. When you give the woman her earrings, she’ll go away, and leave you with a number that you can use to get past the waiting room. Wait a few turns and the door will open.

463. Put the mirror over the access panel.

464. First, move both levers up. The statue will walk into the corridor. Next hint: 425 (Spoiler)

465. Try practicing with the rabbit. After four or five throws, you’ll be pretty good.

466. Wait for him to leave, then pick it up.

467. It’s something you might find on a farm. Next hint: 437 (Spoiler)

468. Move the statue back into the corridor, and you’ll find the entrance where the statue was standing.

469. Fill the bucket with water, then throw the water on the maintenance man.

470. Did you notice the glimmering object? Next hint: 431 (Spoiler)

471. The combination is on a note in the toolbox.

472. Break the mirror; the programmer will leave long enough to look through his things.

473. Open the access panel and go up.

474. A portable radio would have batteries. Next hint: 421 (Spoiler)

475. So, you’ll have to arrange for the group to be out on the porch before you get them to open the door. Next hint: 482 (Spoiler)

476. Put the pillow on the sink before opening the medicine cabinet.

477. You don’t have to, of course, but you’ve now done everything at the house that you need to.

478. Your goal is to join the Hades Explorers’ Society, which is the only way you can move beyond this version of Hell.

479. The controls have been removed, leaving a hole. Did you look there? Next hint: 462 (Spoiler)

480. See the question on the crawlways.

481. How about the rabbit? Next hint: 439 (Spoiler)

482. After getting the ring through the trap door, lead them out onto the porch. Then push the ring down the stairs, and finally lead them back into the house. They’ll open the door and shove the ring down the hall to the den.

483. You can use his absence to look through his desk. Next hint: 428 (Spoiler)

484. It’s the gold skull.

485. You can go west out the hole. You’ll end up in the dumpster.

486. Climb the railing. Next hint: 473 (Spoiler)

487. Read the sign on the gravel path. You’ll notice a cigarette box. Take the cigarette to the teen and offer it to him.

488. You’ll need to wear the parachute prior to the flight, then pull the rip cord while airborne.

489. Have you looked at those animal bones? Next hint: 424 (Spoiler)

Have you tried…

• Visiting the programmer or the elevator operator while wearing the parachute and flippers?

• Participating in the economics experiment after being paid off?

• Praying?

• Looking under the bed in the house in the suburbs?

• Opening the desk drawer while the programmer is in his office?

• Told the priest to “get out”?

• Putting the bedspread on the flaming maintenance man?

• Climbing the trees in the forest?

How the points are scored

13 points for getting off the boat at the start of the game. 50 points for taking each of the amulets (the ivory carving, the gold disk, the jeweled amulet, the silver ring, and the bracelet). 25 points for getting a Celestial Security card. 25 points for getting a Driver’s License. 25 points for getting the 20 Hellmid bill. 13 points for extinguishing the the maintenance man. 13 points for climbing into the rails in the Infernolab lobby for the first time. 25 points for making it up onto the Celestial Security office roof. 13 points for getting the pocket lighter. 25 points for winning at the economics experiment. 13 points for getting rid of the DimNet Lab technician. 13 points for entering Heaven for the first time. 13 points for finding the earrings. 13 points for giving the earrings to the woman in the waiting room. 13 points for getting the RaptureMeal coupon. 13 points for getting the mirror (without breaking it). 13 points for breaking the mirror for the first time. 25 points for getting the programmer’s ID badge. 13 points for getting the DMV appointment card. 13 points for opening the door at the bottom of the crater. 13 points for moving the lion-headed statue for the first time. 13 points for getting rid of the pterodactyl. 13 points for prying open the airplane compartment. 13 points for getting the heat shield. 13 points for getting the flippers. 19 points for getting past the Troll. 13 points for getting in the boat at the end of the game.

You will have to collect all 666 points to complete the game.

Credits

Game design and implementation: Mike Roberts. Testing: Roger Dominick, John Eras, Jeff Laing, Steve McAdams, Bob Newell, Trevor Powell, Lon Thomas. Implemented with TADS: The Text Adventure Development System.